Ahmed has half of a treasure map, which indicates that the treasure is buried in the desert 2x + 6 paces from Castle Rock. Vanessa has the other half, her half indicates that to find the treasure, one must get to Castle Rock, walk x paces to the north, and then walk 2x + 4 paces to the east. If they share their information, then they can find x and save a lot of digging. What is x? Ineed to factor this and use the Pythagorean Theorem for this assignment, please help.
Answer by richwmiller(17219)

2x+6 is the hypotenuse c
x is one leg a
2x+4 is the other leg b
a^2+b^2=c^2
(2x+4)^2+x^2=(2x+6)^2
multiply it out and you get
5 x^2+16 x+16 = 4 x^2+24 x+36
combine terms and you get
x^2-8x-20 = 0
you want factors of -20 which add up to -8
-10 and 2 will do it
(x-10)*(x+2)=0
x=10 and x=-2
discard the negative result.
x=10